OpenStudy (anonymous):

solve for t. t^6+2t^3=63

OpenStudy (anonymous):

Subtract 63 from both sides to get:\[t^6 + 2t^3 - 63 = 0\] Next we'll factor it down to:\[(t^3 + 9) (t^3 -7) = 0\]From there we'll take both sets of parenthesis and set them equal to 0 to get:\[t^3 + 9 = 0\] \[t^3 - 7 = 0\] We need to get t by itself so we subtract the constant number from both sides and cube root each to get: \[t = \sqrt[3]{-9}\] and \[t = \sqrt[3]{7}\]